OpenStudy (alexandrad):

Write the equation of the line that passes through (−2, 6) and (2, 14) in slope-intercept form.

OpenStudy (alexandrad):

A. Y=2X-2 B. Y=2X+10 C. Y=0.5X+7 D. Y=0.4x-4

OpenStudy (steve816):

Do you have a guess?

OpenStudy (alexandrad):

i do not :( i dont know how to do this could you teach me ?

OpenStudy (alexandrad):

is this where you use the formula x1-x2 / y1-y2 ?

OpenStudy (steve816):

OOps! You switched the y and the x. It should be y1-y2 / x1-x2, and yes, you would use the formula.

OpenStudy (alexandrad):

okay thank you again !

OpenStudy (alexandrad):

so i got -8/-4= 2 but what do i do next?

OpenStudy (steve816):

Yay, you got the slope! Now you plug in one of the points to the point slope form! Remember y - y1 = m(x - x1)

OpenStudy (alexandrad):

so i plug in -8 to what

OpenStudy (steve816):

Ok, choose one of your points. Which one would you like to use? You plug in -8 for m.

OpenStudy (alexandrad):

my points as in (-2,6)

OpenStudy (steve816):

Ok, now plug in -2 for x1 and 6 for y1

OpenStudy (steve816):

no no, what you would do is \[y-6=2(x+2)\]

OpenStudy (steve816):

y - 6 = 2x + 4

OpenStudy (steve816):

Answer would be y = 2x + 10

OpenStudy (alexandrad):

:( thank you for your patience lol

OpenStudy (steve816):

No problem alexandra!